OpenStudy (anonymous):

Which group of points lies on the line y = 4x - 3? (1, 1); (2, –1); (–1, –7) (1, 1); (2, 5); (-1, –1) (1, 1); (2, 5); (–1, –7) (1, 1); (2, –1); (–1, –1)

OpenStudy (anonymous):

y=4x-3 (1,1) 1=4*1-3 1=4-3 1=1 (2,-1) (-1)=4*2-3 (-1)=9-3 (-1)=6 (-1,-7) (-7)=4*(-1)-3 (-7)=(-4)-3 (-7)=(-7) (1,1) 1=4*1-3 1=4-3 1=1 (2,5) 5=4*2-3 5=8-3 5=5 (-1,-1) (-1)=4*(-1)-3 (-1)=(-4)-3 (-1)=(-7)